# Gradient Descent for Noisy Optimization

### v4

@@ 2.2 Assumptions
We consider the following two assumptions:
@@
The function $F$ is locally strongly convex.
Remark that Assumption A1 implies that $F$ is strictly convex on $\mathbb{R}^{d}$ and thus that $\theta_{\star}$ is well-defined.
@@
The function $\theta\mapsto f(\theta,z)$ is twice continuously differentiable on $\mathbb{R}^{d}$ for all $z\in\mathbb{R}^{d}$, and
@@
for some $\alpha\in(0,1]$. Moreover, for all (non-empty) compact set $\Theta\subset\mathbb{R}^{d}$ there exists a measurable function $M_{\Theta}:\mathsf{Z}\mapsto[0,\infty)$ such that $\mathbb{E}[M_{\Theta}(Z)]<\infty$ and such that
@@ 2.3 Some implications of Assumptions A1-A2
Assume that A2 holds. Then, for all (non-empty) compact set $\Theta\subset\mathbb{R}^{d}$ there exists a measurable function $M_{\Theta}^{\prime}:\mathsf{Z}\mapsto[0,\infty)$ such that $\mathbb{E}[M_{\Theta}^{\prime}(Z)]<\infty$ and such that, for all $k\in\{0,1,2\}$, we have
@@
By combining Lemma 1 and Lemma 2, it follows that under A2 the function $F$ is locally $L$-smooth. More precisely, under A2, for any (non-empty) compact set $\Theta\subset\mathbb{R}^{d}$ we have
@@
Assume that A2 holds. Then, for all (non-empty) compact set $\Theta\subset\mathbb{R}^{d}$ we have
@@
Assume that A2 holds. Then, for all (non-empty) compact set $\Theta\subset\mathbb{R}^{d}$ there exists a $\theta_{\Theta}\in\Theta$ such that
@@
By combining Lemma 2 and Lemma 4, we readily obtain the following result which notably shows that if $\Theta\subset\mathbb{R}^{d}$ is a (non-empty) compact set then the random function $F_{n}$ is strictly convex on $\Theta$ with $\mathbb{P}$-probability tending to one.
@@
Assume that A1-A2 hold. Then, for all (non-empty) compact set $\Theta\subset\mathbb{R}^{d}$ there exists a constant $c_{\Theta}\in(0,1)$ such that
@@
Assume that A1-A2 hold and let $\Theta\subset\mathbb{R}^{d}$ be a non-empty compact and convex set such that $\theta_{\star}\in\mathring{\Theta}$. Then, $\|\theta_{\star}-\hat{\theta}_{\Theta,n}\|=\mathcal{O}_{\mathbb{P}}(n^{-\frac{\alpha}{1+\alpha}})$, with $\alpha\in(0,1]$ as in A2.
Finally, the following proposition shows that, under Assumptions A1-A2 and in the context of this work (i.e. noisy optimization unded fixed computations budget) the best possible convergence rate for an estimator of $\theta_{\star}$ is $B^{-\frac{\alpha}{1+\alpha}}$, with $\alpha\in(0,1]$ as in A2.
@@ 2.4 Gradient descent with backtracking line search
Algorithm 1 provides a generic description of a gradient descent algorithm that can be used to compute the global minimum $g_{\star}:=\inf_{x\in\mathbb{R}^{d}}g(x)$ of a convex and differentiable function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$. The convergence behaviour of Algorithm 1 depends on the value of the step-size $v_{t}$ chosen on its Line 3 and, in particular, we have the following well-known result (whose proof is given in the appendix for sake of completeness).
@@
Let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a convex and differentiable function such that $g_{\star}>-\infty$. Let $x_{0}\in\mathbb{R}^{d}$ and, for all $T\in\mathbb{N}$, let $x_{T}$ be as computed by Algorithm 1 with $(v_{t})_{t\geq 1}$ such that, for some constant $c>0$,
@@
Then, there exists a $x_{\star}\in\mathbb{R}^{d}$ such that $g(x_{\star})=g_{\star}$ and we have
$g(x_{T})-g_{\star}\leq\frac{\|x_{0}-x_{\star}\|^{2}}{2cT},\quad\forall T\geq 1.$
If the function $g$ is $L$-smooth (assuming without loss of generality that $L>1$) then (2) holds for instance when $v_{t}=v$ for all $t\geq 1$ and some $v\in(0,1/L]$ (see e.g. the proof of Nesterov,, 2018, Theorem 2.1.14, page 80). In practice, the value of $L$ is however rarely known and backtracking line search, described in Algorithm 2, is a simple procedure that can be used to automatically compute, at each iteration $t$, a value for $v_{t}$ which ensures that (2) holds with $c=\beta/L$, where $\beta\in(0,1)$ is a parameter of Algorithm 2.
Interestingly, for the backtracking line search procedure to ensure that (2) holds, it is enough to assume that the function $g$ is locally $L$-smooth, as shown in the next proposition.
@@
Let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a convex, differentiable and locally $L$-smooth function such that $g_{\star}>-\infty$. Let $x_{0}\in\mathbb{R}^{d}$, $\beta\in(0,1)$ and, for all $T\in\mathbb{N}$, let $x_{T}$ be as computed by Algorithm 1 where Algorithm 2 is used to compute $v_{t}$ for all $t\geq 1$. In addition, let $\tilde{K}\subset\mathbb{R}^{d}$ be a compact and convex set such that $\{x\in\mathbb{R}^{d}:\,g(x)\leq g(x_{0})\}\subseteq\tilde{K}$,
$K=\Big{\{}x\in\mathbb{R}^{d}:\,\exists x^{\prime}\in\tilde{K}\text{ such that }\|x-x^{\prime}\|\leq\sup_{\tilde{x}\in\tilde{K}}\|\nabla g(\tilde{x})\|\Big{\}}$
and, noting that $K$ is a compact set, let $L_{K}\in[1,\infty)$ be such that $\|\nabla g(x)-\nabla g(x^{\prime})\|\leq L_{K}\|x-x^{\prime}\|$ for all $x,x^{\prime}\in K$. Then, (2) holds with $c=\beta/L_{K}$.
Given the large literature on convex optimization, the conclusion of Proposition 6 is probably known. We however failed to find a reference establishing a similar result and therefore the proposition is proved in the appendix.
@@ 3.1 A simple application of gradient descent to noisy optimization
In order for the resulting estimator $\tilde{\theta}_{B}$ of $\theta_{\star}$ to be consistent as $B\rightarrow\infty$, it is clear that we must have $n(B)\rightarrow\infty$, so that $F_{n(B)}\rightarrow F$, and $n(B)/B\rightarrow 0$, so that the number of iterations of the algorithm converges to infinity. The following theorem formalizes this intuitive reasoning and provides an upper bound on the convergence rate of $\|\tilde{\theta}_{B}-\theta_{\star}\|$.
@@ 3.2.2 Convergence result
Assuming that A2 holds for $\alpha=\alpha^{\prime}$, it follows from Theorem 2 that taking a small value for $J$ can only make slower the speed at which $\hat{\theta}_{J,B}$ converges to $\theta_{\star}$. However, as mentioned in the introductory section, to observe in practice the $\mathcal{O}\big{(}B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-\delta^{J})}\big{)}$ convergence rate it is needed to minimize the function $F_{n_{J}(B)}$, which may not be the case if $J$ is too large given the available computational budget. For this reason, we expect the observed convergence rate to be $B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-\delta^{J_{B}})}$, with $J_{B}$ the number of calls of Algorithm 3 needed for computing $\hat{\theta}_{J,B}$, that is with $J_{B}$ the smallest value of $j\in\{1,\dots,J\}$ such that $\hat{\theta}_{J_{B},B}=\hat{\theta}_{J,B}$. A sensible approach is therefore to let the algorithm decides how many calls to Algorithm 3 are needed by choosing $J$ large enough so that $J_{B}<J$ with probability close to one. The numerical experiments presented in the next section suggest that $J_{B}$ increases with $B$ at speed $\log B$, and verifying theoretically this observation is left for future research.
@@
Finally, if A2 does not hold for $\alpha=\alpha^{\prime}$ then the proof of Theorem 2 shows that $\|\hat{\theta}_{J,B}-\theta_{\star}\|$ converges to zero at least at speed $B^{-\frac{\alpha}{1+\alpha}(1-\delta)}$, for any $\alpha$ such that A2 holds. Studying more precisely the convergence rate of $\hat{\theta}_{J,B}$ in this context is left for future research, but we expect this rate to be slower than the one obtained when A2 holds for the chosen value of $\alpha^{\prime}$. To understand why, remark that in Algorithm 4 the parameter $\alpha^{\prime}$ only impacts the value of the tolerance parameter used in each call of Algorithm 3, and that the larger $\alpha^{\prime}$ the smaller the tolerance parameter. Then, if $\alpha^{\prime}$ is too large (i.e. if A2 does not hold for $\alpha=\alpha^{\prime}$), for any $j\in\{1,\dots,J\}$ too much computational budget is allocated to the $j$-th call of Algorithm 3, resulting in a waste of computational resources and thus in a slower convergence rate as $B\rightarrow\infty$. In particular, it may be the case that for any values of $B$ we have $J_{B}<J$ with high probability.
@@ 4 Numerical experiments
In this section we study the behaviour of $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$ and of $\mathbb{E}[J_{B}]$ as $B$ increases, as estimated from 100 independent realizations of $(Z_{i})_{i\geq 1}$. In all the experiments presented below, Algorithm 4 is implemented with $\beta=0.5$ and with $C_{\mathrm{grad}}=C_{\mathrm{eval}}=\kappa=\tau=1$, and with $\lceil\kappa B^{\gamma_{j}}\rceil$ replaced by $\max(100,\lceil\kappa B^{\gamma_{j}}\rceil)$ for all $j\geq 1$. It is trivial that Theorem 2 remains valid with this latter modification of the algorithm, which is introduced to prevent to minimize $F_{n_{j}(B)}$ for a value of $n_{j}(B)$ that is too small. In addition, following the discussion of Section 3.2, we let $J=10^{4}$ so that $J_{B}<J$ with probability close to one. Finally, in what follows we denote by $\rho_{J_{B}}$ the coefficient of correlation between the variables $(\mathbb{E}[J_{B}])_{B\geq 1}$ and $(\log B)_{B\geq 1}$, as estimated from 100 independent realizations of $(Z_{i})_{i\geq 1}$. Remark that a value of $\rho_{J_{B}}$ close to one implies that $\mathbb{E}[J_{B}]\approx a+b\log B$ for some constants $a\in\mathbb{R}$ and $b\in(0,\infty)$.

### v5

@@ 2.2 Assumptions
We consider the following two assumptions on the optimization problem (1):
@@
The function $F$ is locally strongly convex, that is, for every compact set $\Theta\subset\mathbb{R}^{d}$ there exists a constant $c_{\Theta}>0$ such that
$F\big{(}a\theta_{1}+(1-a)\theta_{2}\big{)}\leq aF(\theta_{1})+(1-a)F(\theta_{2})-\frac{c_{\Theta}}{2}a(1-a)\|\theta_{1}-\theta_{2}\|^{2},\quad\forall a\in[0,1],\quad\forall(\theta_{1},\theta_{2})\in\Theta^{2}.$
Remark that under A1 we have $F\big{(}a\theta_{1}+(1-a)\theta_{2}\big{)}<aF(\theta_{1})+(1-a)F(\theta_{2})$ for all $a\in[0,1]$ and all $(\theta_{1},\theta_{2})\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, meaning that the function $F$ is strictly convex on $\mathbb{R}^{d}$. Consequently, under A1, the optimization problem (1) is well-defined.
@@
The function $\theta\mapsto f(\theta,z)$ is twice continuously differentiable on $\mathbb{R}^{d}$ for all $z\in\mathsf{Z}$, and
@@
for some $\alpha\in(0,1]$. Moreover, for all compact set $\Theta\subset\mathbb{R}^{d}$ there exists a measurable function $M_{\Theta}:\mathsf{Z}\mapsto[0,\infty)$ such that $\mathbb{E}[M_{\Theta}(Z)]<\infty$ and such that
@@ 2.3 Some implications of Assumptions A1-A2
Assume that A2 holds. Then, for all compact set $\Theta\subset\mathbb{R}^{d}$ there exists a measurable function $M_{\Theta}^{\prime}:\mathsf{Z}\mapsto[0,\infty)$ such that $\mathbb{E}[M_{\Theta}^{\prime}(Z)]<\infty$ and such that, for all $k\in\{0,1,2\}$, we have
@@
By combining Lemma 1 and Lemma 2 it follows that under A2 the function $F$ is locally $L$-smooth. More precisely, under A2, for any compact set $\Theta\subset\mathbb{R}^{d}$ we have
@@
Assume that A2 holds. Then, for all compact set $\Theta\subset\mathbb{R}^{d}$ we have
@@
Assume that A2 holds. Then, for all compact set $\Theta\subset\mathbb{R}^{d}$ there exists a $\theta_{\Theta}\in\Theta$ such that
@@
Under A1 and by combining Lemma 2 and Lemma 4, we readily obtain the following result which notably shows that if $\Theta\subset\mathbb{R}^{d}$ is a compact and convex set then the random function $F_{n}$ is strictly convex on $\Theta$ with $\mathbb{P}$-probability tending to one.
@@
Assume that A1-A2 hold. Then, for all compact set $\Theta\subset\mathbb{R}^{d}$ there exists a constant $c_{\Theta}\in(0,1)$ such that
@@
Assume that A1-A2 hold and let $\Theta\subset\mathbb{R}^{d}$ be a compact and convex set such that $\theta_{\star}\in\mathring{\Theta}$. Then, $\|\theta_{\star}-\hat{\theta}_{\Theta,n}\|=\mathcal{O}_{\mathbb{P}}(n^{-\frac{\alpha}{1+\alpha}})$, with $\alpha\in(0,1]$ as in A2.
Finally, the following proposition shows that, under A1-A2 and in the context of this work (i.e. noisy optimization under fixed computations budget) the best possible convergence rate for an estimator of $\theta_{\star}$ is $B^{-\frac{\alpha}{1+\alpha}}$, with $\alpha\in(0,1]$ as in A2.
@@ 2.4 Gradient descent with backtracking line search
Algorithm 1 provides a generic description of a gradient descent algorithm that can be used to minimize a differentiable function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$. The convergence behaviour of Algorithm 1 depends on the value of the step-size $v_{t}$ chosen on its Line 3, and our theoretical analysis of gradient descent applied to noisy optimization problems builds on the following result.
@@
Let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a differentiable function, $x_{0}\in\mathbb{R}^{d}$, and for all $T\in\mathbb{N}$ let $x_{T}$ be as computed by Algorithm 1 with $(v_{t})_{t\geq 1}$ such that, for some constant $c>0$,
@@
Assume that there exists a convex set $K\subseteq\mathbb{R}^{d}$ such that (i) $x_{t}\in K$ for all $t\geq 0$, (ii) the function $g$ is strictly convex on $K$ and (iii) there exists an $x_{\star}\in K$ such that $g(x_{\star})=\inf_{x\in K}g(x)$. Then,
$g(x_{T})-g(x_{\star})\leq\frac{\|x_{0}-x_{\star}\|^{2}}{2cT},\quad\forall T\geq 1.$
If the function $g$ is $L$-smooth (assuming without loss of generality that $L>1$) then (2) holds for instance when $v_{t}=v$ for all $t\geq 1$ and some $v\in(0,1/L]$ (see e.g. the proof of Nesterov,, 2018, Theorem 2.1.14, page 80). In practice, the value of $L$ is however rarely known and backtracking line search, described in Algorithm 2, is a simple procedure that can be used to automatically compute, at each iteration $t$, a value for $v_{t}$ which ensures that (2) holds with $c=\beta/L$, where $\beta\in(0,1)$ is a parameter of Algorithm 2. This property of backtracking line search is formalized in the following proposition, which plays a key role in the proofs of our main results.
@@
Let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a differentiable function, $x_{0}\in\mathbb{R}^{d}$, $\beta\in(0,1)$, and for all $T\in\mathbb{N}$ let $x_{T}$ be as computed by Algorithm 1 where Algorithm 2 is used to compute $v_{t}$ for all $t\geq 1$. In addition, let $\tilde{K}\subseteq\mathbb{R}^{d}$ be a convex set containing the set $\{x\in\mathbb{R}^{d}:\,g(x)\leq g(x_{0})\}$ and let
$K=\Big{\{}x\in\mathbb{R}^{d}:\,\exists x^{\prime}\in\tilde{K}\text{ such that }\|x-x^{\prime}\|\leq\sup_{\tilde{x}\in\tilde{K}}\|\nabla g(\tilde{x})\|\Big{\}}.$
Assume that the function $g$ is strictly convex on $K$ and that there exists a constant $L_{K}\in[1,\infty)$ such that $\|\nabla g(x)-\nabla g(x^{\prime})\|\leq L_{K}\|x-x^{\prime}\|$ for all $(x,x^{\prime})\in K^{2}$. Then, (2) holds with $c=\beta/L_{K}$.
Given the large literature on convex optimization, the conclusions of Propositions 5-6 are probably known. We however failed to find a reference establishing similar results and therefore the two propositions are proved in the appendix.
@@ 3.1 A simple application of gradient descent to noisy optimization
In order for the resulting estimator $\tilde{\theta}_{B}$ of $\theta_{\star}$ to be consistent as $B\rightarrow\infty$, it is clear that we must have (i) $n(B)\rightarrow\infty$, so that $F_{n(B)}\rightarrow F$, and (ii) $n(B)/B\rightarrow 0$, so that the number of iterations of the algorithm converges to infinity. The following theorem formalizes this intuitive reasoning and provides an upper bound on the convergence rate of $\|\tilde{\theta}_{B}-\theta_{\star}\|$.
@@ 3.2.2 Convergence result
Assuming that A2 holds for $\alpha=\alpha^{\prime}$, it follows from Theorem 2 that taking a small value for $J$ can only make slower the speed at which $\hat{\theta}_{J,B}$ converges to $\theta_{\star}$. However, as mentioned in the introductory section, to observe in practice the $\mathcal{O}\big{(}B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-\delta^{J})}\big{)}$ convergence rate it is needed to minimize the function $F_{n_{J}(B)}$, which may not be the case if $J$ is too large given the available computational budget. For this reason, we expect the observed convergence rate to be $B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-\delta^{J_{B}})}$, with $J_{B}$ the number of calls of Algorithm 3 needed for computing $\hat{\theta}_{J,B}$, that is with $J_{B}$ the smallest value of $j\in\{1,\dots,J\}$ such that $\hat{\theta}_{J_{B},B}=\hat{\theta}_{J,B}$. A sensible approach is therefore to let the algorithm decides how many calls to Algorithm 3 are needed by choosing $J$ large enough so that $J_{B}<J$ with probability close to one. The numerical experiments presented in the next section suggest that $J_{B}$ increases with $B$ at speed $\log(B)$, and verifying theoretically this observation is left for future research.
@@
Finally, if A2 does not hold for $\alpha=\alpha^{\prime}$ then the proof of Theorem 2 shows that $\|\hat{\theta}_{J,B}-\theta_{\star}\|$ converges to zero at least at speed $B^{-\frac{\alpha}{1+\alpha}(1-\delta)}$, for any $\alpha$ such that A2 holds. Studying more precisely the convergence rate of $\hat{\theta}_{J,B}$ in this context is left for future research, but we expect this rate to be slower than the one obtained when A2 holds for the chosen value of $\alpha^{\prime}$. To understand why, remark that in Algorithm 4 the parameter $\alpha^{\prime}$ only impacts the value of the tolerance parameter used in each call of Algorithm 3, and that the larger $\alpha^{\prime}$ the smaller the tolerance parameter. Then, if $\alpha^{\prime}$ is too large (i.e. if A2 does not hold for $\alpha=\alpha^{\prime}$), for any $j\in\{1,\dots,J\}$ too much computational budget is allocated to the $j$-th call of Algorithm 3, resulting in a waste of computational resources and thus in a slower convergence rate as $B\rightarrow\infty$. In particular, it may be the case that even for large values of $B$ we have $J_{B}<J$ with high probability.
@@ 3.2.4 Understanding the convergence rate given in Theorem 2
3.3 Validity of GD-BLS for solving noisy optimization problems: sketch of proof
We end this section with a sketch of proof of the validity of GD-BLS for solving noisy optimization problems, the goal being to provide the reader with some insights about how Assumptions A1-A2 and the properties of the GD-BLS algorithm are used to prove our main results. Specifically, in this subsection we let $\theta_{0}\in\mathbb{R}^{d}$ and consider the problem of solving the optimization problem (1) by applying the GD-BLS algorithm (i.e. Algorithms 1-2) on the random function $F_{n}$, using $\theta_{0}$ as starting value. We denote by $(\theta^{(n)}_{t})_{t\geq 1}$ the resulting sequence of iterates and below we show, in an informal way, that the results of Sections 2.3-2.4 imply that $\|\lim_{t\rightarrow\infty}\theta_{t}^{(n)}-\theta_{\star}\|={\scriptscriptstyle\mathcal{O}}_{\mathbb{P}}(1)$.
To this aim remark first that, by using the descent property of GD-BLS, all the elements of $(\theta^{(n)}_{t})_{t\geq 1}$ belong to the (random) set $\tilde{\Theta}_{0,n}:=\{\theta\in\mathbb{R}^{d}:\,F_{n}(\theta)\leq F_{n}(\theta_{0})\}$. Next, by using Lemma 3 and under A1, it can be proved that there exists a compact and convex set $\tilde{\Theta}_{0}$ such that $\tilde{\Theta}_{0,n}\subset\tilde{\Theta}_{0}$ with $\mathbb{P}$-probability tending to one. To proceed further we let
$\Theta_{0,n}=\Big{\{}\theta\in\mathbb{R}^{d}:\,\exists\theta^{\prime}\in\tilde{\Theta}_{0}\text{ such that }\|\theta-\theta^{\prime}\|\leq\sup_{\tilde{\theta}\in\tilde{\Theta}_{0}}\|\nabla F_{n}(\tilde{\theta})\|\Big{\}}$
$\Theta_{0}=\Big{\{}\theta\in\mathbb{R}^{d}:\,\exists\theta^{\prime}\in\tilde{\Theta}_{0}\text{ such that }\|\theta-\theta^{\prime}\|\leq\sup_{\tilde{\theta}\in\tilde{\Theta}_{0}}\|\nabla F(\tilde{\theta})\|+1\Big{\}}$
and remark that, by Lemma 3, with $\mathbb{P}$-probability tending to one we have $\Theta_{0,n}\subset\Theta_{0}$. Under A1 the set $\Theta_{0}$ is compact and convex and thus, by Lemma 1 and Corollary 1, it follows that with $\mathbb{P}$-probability tending to one the function $F_{n}$ is strictly convex and $L$-smooth on $\Theta_{0}$ (with $L\in(0,\infty)$ a finite constant). Without loss of generality we can assume that $\Theta_{0}$ is sufficiently large so that $\theta_{\star}$ belongs to the interior of this set, in which case $\|\hat{\theta}_{\Theta_{0},n}-\theta_{\star}\|={\scriptscriptstyle\mathcal{O}}_{\mathbb{P}}(1)$ by Lemma 5. To sum-up, with $\mathbb{P}$-probability tending to one (i) the function $F_{n}$ is strictly convex and $L$-smooth on the compact and convex set $\Theta_{0}$, (ii) $F_{n}(\hat{\theta}_{\Theta_{0},n})=\inf_{\theta\in\Theta_{0}}F_{n}(\theta)$ and (iii) the sequence of iterates $(\theta^{(n)}_{t})_{t\geq 1}$ generated by the GD-BLS algorithm applied on $F_{n}$ is a sequence in $\tilde{\Theta}_{0}$. Then, by using Propositions 5-6, it readily follows that $\|\lim_{t\rightarrow\infty}\theta_{t}^{(n)}-\hat{\theta}_{\Theta_{0},n}\|={\scriptscriptstyle\mathcal{O}}_{\mathbb{P}}(1)$, and since $\|\hat{\theta}_{\Theta_{0},n}-\theta_{\star}\|={\scriptscriptstyle\mathcal{O}}_{\mathbb{P}}(1)$ we can conclude that $\|\lim_{t\rightarrow\infty}\theta_{t}^{(n)}-\theta_{\star}\|={\scriptscriptstyle\mathcal{O}}_{\mathbb{P}}(1)$.
@@ 4 Numerical experiments
In this section we study the behaviour of $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$ and of $\mathbb{E}[J_{B}]$ as $B$ increases, as estimated from 100 independent realizations of $(Z_{i})_{i\geq 1}$. In all the experiments presented below, Algorithm 4 is implemented with $\beta=0.5$ and with $C_{\mathrm{grad}}=C_{\mathrm{eval}}=\kappa=\tau=1$, and with $\lceil\kappa B^{\gamma_{j}}\rceil$ replaced by $\max(100,\lceil\kappa B^{\gamma_{j}}\rceil)$ for all $j\geq 1$. It is trivial that Theorem 2 remains valid with this modification of the algorithm, which is introduced to prevent to minimize $F_{n_{j}(B)}$ for a value of $n_{j}(B)$ that is too small. In addition, following the discussion of Section 3.2, we let $J=10^{4}$ so that $J_{B}<J$ with probability close to one. Finally, in what follows we denote by $\rho_{J_{B}}$ the coefficient of correlation between the variables $(\mathbb{E}[J_{B}])_{B\geq 1}$ and $(\log(B))_{B\geq 1}$, as estimated from 100 independent realizations of $(Z_{i})_{i\geq 1}$. Remark that a value of $\rho_{J_{B}}$ close to one implies that $\mathbb{E}[J_{B}]\approx a+b\log(B)$ for some constants $a\in\mathbb{R}$ and $b\in(0,\infty)$.